TSS problem confirmed again.

Test disable/enable via sw: no patch. The whistle come again.

saved data:

~/adopt/measures/adsec_calib/status_save/datifischio_26052012.sav

System put upside down.

As first, all the TSS connectors/levels were inspected and the level were reduced to the LBT one dismuounting each crate.

On crate 4 found the DC DC converter con le viti spanate, quindi non e' stato possibile cambiarlo

All cleaned.

Put mirror back in vertical. Found actuator con bad readings (? poor value ?) DSP #273 mirror. (6um costanti, contro 15 micron dei vicini

Premuto un po' e' passato al bianco e poi e' rinvivito

Test 2xset shell -> tutto ok; Test TSS a shell rippata: -> tutto ok: /home/aosup/adopt/measures/adsec_calib/spi/20120527_032238/spi_after_tss_fix.sav

Shell a 25 gradi: 20 min storia temporale: ->misura spi a 5 min: ~/adopt/measures/adsec_calib/spi/20120527_040339 Set and RIP: spi: ~/adopt/measures/adsec_calib/spi/20120527_041006

Set and RIP: SPI ~/adopt/measures/adsec_calib/spi/20120527_041913

All OK! -- MarcoXompero - 26 May 2012
Topic revision: r4 - 27 May 2012, MarcoXompero
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback